# DUNIA: Pixel-Sized Embeddings via Cross-Modal Alignment for Earth Observation Applications

### v1

@@ 4.2.2 Finetuning Performance
Finetuning results in Table 2 demonstrate that DUNIA outperforms the other models in vertical structure-related products, surpassing both AnySat and CROMA, with CROMA trailing significantly. In comparison to DUNIA’s zero-shot results, all the models underperformed in the finetuning setting for $\mathcal{W}_{c}$ and $\mathcal{W}_{pai}$ due to the long-tailed distribution of these two products that are not well modeled with the simple $L_{1}$ loss that we used. Qualitatively, as shown in Figure 3, DUNIA and AnySat show a similar level of detail; however, DUNIA is capable of producing higher values, which we attribute to the alignment with the vertical structure in the pre-training stage. For CROMA, Figure 3 shows that this model produces much smoother maps than the two other models.
@@
For $CLC_{+}$ and $PF$, DUNIA’s performance in the fine-tuning setting increased in comparison to the zero-shot setting with results on par with the much more data-heavy AnySat model (Table 2). However, DUNIA underperformed by a significant margin ($\downarrow 4.1\%$) in comparison to AnySat for the $PASTIS$ dataset. Qualitatively (Figure 3) both DUNIA and AnySat show similar level of detail for the $CLC_{+}$ dataset, in contrast to CROMA which also showed very smooth maps with fewer details.
@@
In the novel task of waveform generation, we observed a correlation ($r$) increase of .08 between reference and generated waveforms compared to retrieval ($r$=.70 with S=50K l, KNN=1). This correlation dropped to .75 when the diffusion model was trained on just 20% of the available waveforms. Retrieved vs. generated waveforms can be found in Figures 5, 6, 7 and 8.
@@
To further analyze the contributions of different architectural choices, including the use of separate decoders and the impact of our alignment strategies, we provide additional ablation studies in Appendix E
@@ 5 Discussion
We introduced DUNIA, a novel framework for Earth Observation applications that learns dense representations through mono-modal and cross-modal alignment. This strategy enables our model to perform mono-modal and cross-modal retrieval tasks on several datasets with very high precision. It also allows novel capabilities like retrieving or generating highly complex outputs from pixel inputs. To the best of our knowledge, this is the only model capable of generating GEDI waveforms at this scale.
Our main contributions lie in several design choices at the decoding stage. The proposed pre-trained model encoder is flexible and can be adapted to components from other existing approaches; for instance, modules that support multi-temporal input, or handle images with varying resolutions. Nonetheless, in this work we focused on input data - Sentinel- 1 &2, alongside GEDI - that are freely accessible globally to promote easier adoption.
One of the main limitations of this work is also coincidentally the reliance on static inputs from Sentinel- 1& 2 data, which has limitations for datasets requiring multi-temporal images, such as crop type mapping. However, our design choices prioritize mono-temporal data to reduce storage requirements and enable applications in areas where time series data are not readily available. Another limitation, common to all EO models, is that the current pre-trained model is tailored to the area and year on which it was trained on; we expect that it will require pre-training when used elsewhere. Nonetheless, the model was designed to not be computationally expensive and can be performed with limited computational resources.
@@
Current forest monitoring systems heavily rely on ground-based inventories, which provide robust statistical estimates of biomass and carbon at national or regional scales but lack the spatial granularity needed for tracking localized carbon gains and losses. Additionally, the largest forested regions—particularly in boreal and tropical areas—remain vastly under-sampled. While satellite data offers global coverage, traditional remote sensing approaches are often constrained by their reliance on extensive labeled datasets, limiting their ability to map forest properties comprehensively.
DUNIA introduces a novel self-supervised learning approach that overcomes these limitations by generating pixel-level embeddings from freely accessible satellite data, including optical and radar imagery. By aligning these embeddings with sparse but highly informative LiDAR waveforms, our approach enables the estimation of multiple forest attributes—including canopy height, fractional cover, land cover, tree species, and vertical structure—without requiring task-specific labels. This method allows for zero-shot and low-shot predictions of key forest variables, paving the way for consistent and scalable monitoring of global ecosystems.
@@ Appendix A Implementation Details
A.1 Pre-training Model
@@
To reconstruct high-resolution embeddings, feature maps from deeper layers ($d=4$) are progressively upsampled using bilinear interpolation and refined through two $3\times 3$ Conv-BatchNorm-ReLU operations before being concatenated with corresponding features from earlier transformer layers. A final $1\times 1$ convolution ensures that the final output dimension matches $D_{p}$ when required.
This architecture is instantiated as two separate decoders, outputting $\mathcal{O}^{\mathcal{V}}$ and $\mathcal{O}^{\mathcal{H}}$ at the highest resolution.
@@ A.3 Multitemporal Image Processing Model
A.4 Waveform Generation
The corresponding objective of a DM can then be written as:
where the target $y$ can be the input noise $\epsilon$, the original input $x$ variable, or the velocity $v=\alpha_{t}\epsilon-\sigma_{t}x$. DMs can also be parametrized by a condition $c\in\mathbb{R}^{D}$ in addition to the diffusion time $t$, and the corresponding objective becomes:
Appendix B Experimental Settings
Sentinel-2 Data.
We used Level-2A surface reflectance data (S2-L2A) from Google Earth Engine (GEE). The dataset includes bands at 10m (Blue, Green, Red, and NIR) and 20m (Red Edge 1-4, SWIR1, and SWIR2) spatial resolutions, all upscaled to 10m ground sampling distance (GSD) using cubic interpolation. We created two sets of mosaics: a single mosaic during the leaf-on season using the median of all available images from April to September 2020. This mosaic is used as input for the pre-trained model. The other set is used as input for the multitemporal AE and is comprised of three mosaics, each representing the median acquisitions of all images in a four-month span from October 2019 until September 2020. Cloudy pixels were filtered out using the S2 Cloud Probability dataset provided by SentinelHub in GEE.
Sentinel-1 Data.
Sentinel-1 (S1) data were obtained from the Sentinel-1A and Sentinel-1B satellites, operating at the C-band ( 6 cm wavelength). Images were collected in Interferometric Wide swath mode (IW) with VV and VH polarizations, derived from the high-resolution Level-1 ground range detected (GRD) product. The original 20m × 22m resolution was resampled to 10m × 10m GSD. Similar to S2, we also created two sets of mosaics. The S1 data were calibrated using the Sentinel SNAP toolbox, converting pixel values to backscattering coefficients ($\sigma^{0}$) in linear units and applying geometric correction using the 30m Shuttle Radar Topography Mission (SRTM) Digital Elevation Model (DEM). Finally, the backscattering coefficients $\sigma_{\theta}^{0}$ acquired at difference incidence angles $\theta$ were normalized to a common reference incidence angle set at $\sigma_{ref}=40^{\circ}$ using the cosine correction equation (N. Baghdadi & Neeson (2001); Topouzelis et al. (2016)):
GEDI Data.
[…]
PureForest ($PF$).
CLC+Backbone ($CLS_{+}$).
PASTIS dataset.
Vertical Structure dataset.
Appendix C Objective Functions.
@@ Appendix B Experimental Settings
Appendix D Model Configuration and Optimization
D.1 Model Configuration
The image model is pre-trained on $64\times 64$ pixel sized images with 14 bands, a 16-layer transformer with 8 attention heads, a patch size of eight, 512 embedding dimensions, a feedforward layer with two linear layers with a hidden dimension of 2048, GeGLU activation, and 0.1 dropout and attention dropout rates. The two image decoders are identical and project the decoded image feature map to 64 embedding dimensions (i.e. $D_{p}=64$). Finally, we set the window size $w$ of the NA layers to 19 pixels.
The waveform AE is trained on 256 bin-long waveforms, with RVQ layers composed of eight quantizers, and each quantizer containing 512 codebook entries. As such, the waveform encoder encodes the waveforms into a $16\times 16$ latent representation which is average pooled along the channel dimension and projected to 64 embeddings dimension for the pixel-waveform alignment.
The multi-temporal image AE is trained on three $64\times 64$ pixel sized input images, each with 14 bands, and produces a similarly shaped feature map on the temporal and spatial dimensions, and 64 embeddings dimension. This output is then average-pooled on the temporal dimension for the pixel-pixel alignment.
Finally, the diffusion model is trained on the $16\times 16$ quantized waveform latents, which are first projected to 128 channels before being passed to the UNet model.
D.2 Training, Inference, and Generation
@@
Appendix E Ablation Study
E.1 Contrastive loss choice
@@
We used two different loss functions for the Pixel-Pixel alignment and Pixel-Waveform alignment. Table 3 shows that the cosine similarity (CS) using the VICReg loss for the pixel-waveform alignment only reaches a maximum of 0.56 for the positive pairs and 0.10 for the negative pairs, indicating poor alignment. On the other hand, using the ZERO-CL loss, the CS between the positive pairs reaches 0.86, and 0.35 between the negative pairs. We attribute the poor results for Pixel-Waveform alignment using the VICReg loss to the low number of available waveforms within a mini-batch. In contrast, for the Pixel-Pixel alignment, the CS for the positive pairs is 0.99, and 0.04 for the negative pairs using the VICReg loss. It decreases to 0.98 for the positive pairs and 0.45 for the negative pairs with ZERO-CL. Nevertheless, using ZERO-CL for both modalities increased training times exponentially due to the whitening transformation and the high number of available pixels within the mini-batch.
@@
Table 3: Cosine similarity between positive pairs (+) and negative pairs (-) during training. P represents a pixel embedding, W represents a waveform embedding. Positive Pairs consist of a pixel embedding and its corresponding pixel/waveform embedding, while negative pairs are computed between a pixel embedding and other pixel/waveform embeddings within the mini-batch.
@@
E.2 Using a Shared Decoder
In our formulation, we used two different decoders for horizontal and vertical structural understanding. We hypothesized that a single real-valued embedding cannot simultaneously encode contrasting information, such as trees of the same species with different vertical structures. Table 4 Shows that in the retrieval case, training a single decoder produces pixel embeddings with no semantic overlap with the waveforms with an average CS of -0.42, while the retrieved pixel embeddings given a pixel input are still highly similar. In the case of training separate decoders, retrieval performance is drastically different, with high correlations (0.99) between pixels and the retrieved waveforms.

### v2

@@ 4.2.2 Fine-tuning Performance
Fine-tuning results in Table 2 demonstrate that DUNIA outperforms the other models in vertical structure-related products, with all models except AnySat trailing significantly. In comparison to DUNIA’s zero-shot results, all the models underperformed in the finetuning setting for $\mathcal{W}_{c}$ and $\mathcal{W}_{pai}$ due to the long-tailed distribution of these two products that are not well modeled with the simple $L_{1}$ loss that we used. Qualitatively, as shown in Figures I1 and I2, DUNIA and AnySat show a similar level of detail; however, DUNIA is capable of producing higher values, which we attribute to the alignment with the vertical structure in the pre-training stage. For CROMA, Figures I1 and I2 shows that this model produces much smoother maps than the two other models.
@@
For $CLC_{+}$ and $PF$, DUNIA’s performance in the fine-tuned setting increased in comparison to the zero-shot setting with results on par with the much more data-heavy AnySat model (Table 2). However, DUNIA underperformed by a significant margin ($\downarrow 4.1\%$) in comparison to AnySat for the $PASTIS$ dataset. Qualitatively (Figure I3) both DUNIA and AnySat show similar level of detail for the $CLC_{+}$ dataset, in contrast to CROMA which also showed very smooth maps with fewer details.
@@
In the novel task of waveform generation, we observed a correlation ($r$) increase of .08 between reference and generated waveforms compared to retrieval ($r$=.70 with S=50K l, KNN=1). This correlation dropped to .75 when the diffusion model was trained on just 20% of the available waveforms (Table G2). Retrieved vs. generated waveforms can be found in Figures I5, I6, I7 and I8.
@@
To further analyze the contributions of different architectural choices, including the use of separate decoders and the impact of our alignment strategies, we provide additional ablation studies in Appendix H.
@@ 5 Discussion
We introduced DUNIA, a novel framework for Earth Observation applications that learns dense representations through mono-modal and cross-modal alignment. This strategy enables our model to perform mono-modal and cross-modal retrieval tasks on several datasets with high precision. It also allows novel capabilities like retrieving or generating highly complex outputs from pixel inputs. To the best of our knowledge, this is the only model capable of generating GEDI waveforms at this scale.
Our main contributions lie in several design choices at the decoding stage. The proposed pre-trained model encoder is flexible and can be adapted to components from other existing approaches; for instance, modules that support multi-temporal input, or handle images with varying resolutions. Nonetheless, in this work we focused on input data - S-1 & 2, alongside GEDI - that are freely accessible globally to promote easier adoption.
One of the main limitations of this work is also coincidentally the reliance on static inputs from S-1 & 2 data, which has limitations for datasets requiring multi-temporal images, such as crop type mapping. However, our design choices prioritize mono-temporal data to reduce storage requirements and enable applications in areas where time series data are not readily available. Another limitation, common to all EO models, is that the current pre-trained model is tailored to the area and year on which it was trained on; we expect that it will require pre-training when used elsewhere. Nonetheless, the model was designed to be pre-trained with limited computational resources.
Acknowledgements
This work is supported via the AI4Forest project, which is funded by the French National Research Agency (ANR; grant number ANR-22-FAI1-0002) and the German Federal Ministry of Education and Research (BMBF; grant number 01IS23025A).
@@ Acknowledgements
Current forest monitoring systems heavily rely on ground-based inventories, which provide robust statistical estimates of biomass and carbon at national or regional scales but lack the spatial granularity needed for tracking localized carbon gains and losses. Additionally, the largest forested regions—particularly in boreal and tropical areas—remain vastly under-sampled. While satellite data offers global coverage, traditional approaches are often constrained by their reliance on extensive labeled datasets, limiting their ability to map forest properties comprehensively.
DUNIA introduces a novel self-supervised learning approach that overcomes these limitations by generating pixel-level embeddings from freely accessible satellite data, including optical and radar imagery. By aligning these embeddings with sparse but highly informative LiDAR waveforms, our approach enables the estimation of multiple forest attributes—including canopy height, fractional cover, land cover, tree species, and vertical structure—without requiring task-specific labels. This method allows for zero-shot and low-shot predictions of key forest variables, paving the way for scalable monitoring of global ecosystems.
@@ Appendix A Implementation Details
In the main text, we provided a high-level overview of the architecture and training setup. Here, we include additional implementation details for key modules and components of our framework. These details were omitted from the main body for brevity and are provided here for completeness and reproducibility.
Figure A1: An overview of our framework showcasing the main blocks and connections. The pre-trained model, which take as input a single date S-1 & 2 image mosaic, is shaded in blue. The multi-temporal image AE, which take as input a multi-date S-1& 2 image mosaic, is shaded in yellow. The vertical structure AE, which take as input the GEDI waveforms, is shaded in green.
A.1 Pre-Trained Model
@@
To reconstruct high-resolution embeddings, feature maps from deeper layers ($d=4$) are progressively upsampled using bilinear interpolation and refined through two $3\times 3$ Conv-BatchNorm-ReLU operations before being concatenated with corresponding features from earlier transformer layers. A final $1\times 1$ convolution ensures that the final output dimension matches $D_{p}$ when required. This architecture is instantiated as two separate decoders, outputting $\mathcal{O}^{\mathcal{V}}$ and $\mathcal{O}^{\mathcal{H}}$.
@@ A.3 Multitemporal Image Processing Model
Appendix B Objective Functions.
@@
Appendix C Waveform Generation
The corresponding objective of a DM can then be written as:
where $p_{data}$ denotes the data distribution over the clean inputs $x_{0}$, and the target $y$ can be the input noise $\epsilon$, the original input $x$ variable, or the velocity $v=\alpha_{t}\epsilon-\sigma_{t}x$. DMs can also be parametrized by a condition $c\in\mathbb{R}^{D}$ in addition to the diffusion time $t$, and the corresponding objective becomes:
Appendix D Visual Comparison: Pixel vs. Patch Embeddings
To highlight the difference in spatial granularity between our proposed pixel-level embedding model and a patch-based baseline (e.g., CROMA), we present a side-by-side qualitative visualization in Figure D1. Both models are applied to the same S-1 & 2 input, and their embeddings are projected to RGB using t-SNE. The pixel-based model captures finer spatial structures and preserves object boundaries more effectively than the patch-based baseline, yielding lower-resolution outputs.
Figure D1: Comparison between pixel-level embeddings from our proposed model (top row) and patch-based embeddings from one of the baseline models (bottom row), both generated from the same S-1 & 2 inputs (left). Embeddings are visualized using t-SNE.
Appendix E Model Configuration and Optimization
E.1 Model Configuration
The image model is pre-trained on $64\times 64$ pixel sized images with 14 bands, a 16-layer transformer with 8 attention heads, a patch size of eight, 512 embedding dimensions, a feedforward layer with two linear layers with a hidden dimension of 2048, GeGLU activation, and 0.1 dropout and attention dropout rates. The two image decoders are identical and project the decoded image feature map to 64 embedding dimensions (i.e. $D_{p}=64$). Finally, we set the window size $w$ of the NA layers to 19 pixels. The waveform AE is trained on 256 bin-long waveforms, with RVQ layers composed of eight quantizers, and each quantizer containing 512 codebook entries. As such, the waveform encoder encodes the waveforms into a $16\times 16$ latent representation which is average pooled along the channel dimension and projected to 64 embeddings dimension for the pixel-waveform alignment. The multi-temporal image AE is trained on three $64\times 64$ pixel sized input images, each with 14 bands, and produces a similarly shaped feature map on the temporal and spatial dimensions, and 64 embeddings dimension. This output is then average-pooled on the temporal dimension for the pixel-pixel alignment. Finally, the diffusion model is trained on the $16\times 16$ quantized waveform latents, which are first projected to 128 channels before being passed to the UNet model.
E.2 Training, Inference, and Generation
@@
Appendix F Experimental Settings
Sentinel-2 Data.
We used Level-2A surface reflectance data (S2-L2A) from Google Earth Engine (GEE). The dataset includes bands at 10m (Blue, Green, Red, and NIR) and 20m (Red Edge 1-4, SWIR1, and SWIR2) spatial resolutions, all upscaled to 10m ground sampling distance (GSD) using cubic interpolation. We created two sets of mosaics: a single mosaic during the leaf-on season using the median of all available images from April to September 2020. This mosaic is used as input for the pre-trained model. The other set is used as input for the multitemporal AE and is comprised of three mosaics, each representing the median acquisitions of all images in a four-month span from October 2019 until September 2020. Cloudy pixels were filtered out using the S-2 Cloud Probability dataset provided by SentinelHub in GEE.
Sentinel-1 Data.
Sentinel-1 data were obtained from the Sentinel-1A and Sentinel-1B satellites, operating at the C-band ( 6 cm wavelength). Images were collected in Interferometric Wide swath mode (IW) with VV and VH polarizations, derived from the high-resolution Level-1 ground range detected (GRD) product. The original 20 m × 22 m resolution was resampled to 10 m × 10 m GSD. Similar to S-2, we also created two sets of mosaics. The S-1 data were calibrated using the Sentinel SNAP toolbox, converting pixel values to backscattering coefficients ($\sigma^{0}$) in linear units and applying geometric correction using the 30m Shuttle Radar Topography Mission (SRTM) Digital Elevation Model (DEM). Finally, the backscattering coefficients $\sigma_{\theta}^{0}$ acquired at difference incidence angles $\theta$ were normalized to a common reference incidence angle set at $\sigma_{ref}=40^{\circ}$ using the cosine correction equation (N. Baghdadi & Neeson (2001); Topouzelis et al. (2016)):
GEDI Data.
GEDI measures the vertical structure of objects (e.g., vegetation) using three lasers emitting near-infrared light (1064 nm), one of the laser beams is split in two, and after applying optical dithering, this results in eight ground tracks spaced 600 m apart, with each shot having a 25 m footprint. For this study, we used GEDI Level 1B (L1B), Level 2A (L2A), and Level 2B (L2B) data from April 2019 to December 2021. From the L1B product we extracted the waveforms and their geolocation (i.e., latitude, longitude, and elevation), and the geolocation of the GEDI instrument for each shot. Finally, as the waveforms are stored as vectors of 1420 bins (1 bin = 0.15 m), we also extracted for each waveform the bin count, the signal start ($\mathcal{W}_{start}$) and signal end ($\mathcal{W}_{end}$). The latter two indicate the location of the canopy top and the end position of the waveform before the noise, respectively. From the L2A data product, we extracted the $RH_{98}$ (referred to $\mathcal{W}_{rh}$), which represents the estimated height of the tallest object within the waveform footprint, the canopy cover ($\mathcal{W}_{c}$). Finally, from the L2B data product, we extracted the Plant Area Index ($\mathcal{W}_{pai}$).
PureForest ($PF$).
CLC+Backbone ($CLS_{+}$).
PASTIS dataset.
Vertical Structure dataset.
Appendix G Impact of Label Quantity on Performance
Table G1: Top-1 retrieval-based zero-shot classification performance of DUNIA for different database sizes. and are DUNIA query embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples in the retrieval database, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. $\text{KNN}_{b}$ represents the best KNN value for a given dataset. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m.
Dataset Metric S $\text{KNN}_{b}$ 100% S 10% S 5% S
$\mathcal{W}_{rh}$ RMSE (r) 50K l 50 2.0 (.93) 2.1 (.92) 2.1 (.92)
$\mathcal{W}_{c}$ RMSE (r) 50K l 50 11.7 (.89) 12.4 (.86) 12.0 (.84)
$\mathcal{W}_{pai}$ RMSE (r) 50K l 50 0.71 (.75) 0.72 (.75) 0.72 (.74)
$CLC_{+}$ wF1 500 im 50 80.1 75.2 70.2
$PASTIS$ OA 500 im 50 56.2 52.4 48.3
$PF$ wF1 50K l 5 76.0 73.5 70.9
$\mathcal{W}^{**}$ r 50K l 1 .70 .67 .66
Table G2: Fine-tuning performance of DUNIA and the five competing models. and are DUNIA’s embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively. S represents the number of samples used for the fine-tuning, with im meaning a $64\times 64$ pixels fully annotated image, and l meaning a single annotated pixel. $\mathcal{W}^{**}$ represents performance results for vertical structures higher than 5 m. Best scores are in bold.
Dataset Metric Samples (S) DUNIA AnySat Croma DOFA DeCUR SatMAE
$\mathcal{W}_{rh}$ RMSE (r) 1K im 1.4 (.93) 2.8 (.89) 3.6 (.76) 11.2 (.50) 11.1 (.52) 10.5 (.52)
$\mathcal{W}_{c}$ RMSE (r) 1K im 10.0 (.83) 12.4 (.79) 14.5 (.72) 30.1 (.48) 29.4 (.47) 30.7 (.47)
$\mathcal{W}_{pai}$ RMSE (r) 1K im 0.61 (.70) 0.94 (.67) 1.5 (.60) 1.7 (.35) 1.7 (.36) 1.9 (.37)
$CLC_{+}$ wF1 1K im 89.4 89.5 85.9 71.1 74.6 73.8
$PASTIS$ wF1 300K im 75.3 80.2 71.0 52.2 56.4 54.8
$PF$ wF1 10K l 80.1 80.0 80.2 79.8 78.5 78.7
$\mathcal{W}^{**}$ r $\approx$3.8M l .75
Appendix H Ablation Studies
H.1 Loss Choice
@@
We used two different loss functions for the Pixel-Pixel alignment and Pixel-Waveform alignment. Table H1 shows that the cosine similarity (CS) using the VICReg loss for the pixel-waveform alignment only reaches a maximum of 0.56 for the positive pairs and 0.10 for the negative pairs, indicating poor alignment. On the other hand, using the ZERO-CL loss, the CS between the positive pairs reaches 0.86, and 0.35 between the negative pairs. We attribute the poor results for Pixel-Waveform alignment using the VICReg loss to the low number of available waveforms within a mini-batch. In contrast, for the Pixel-Pixel alignment, the CS for the positive pairs is 0.99, and 0.04 for the negative pairs using the VICReg loss. It decreases to 0.98 for the positive pairs and 0.45 for the negative pairs with ZERO-CL. Nevertheless, using ZERO-CL for both modalities increased training times exponentially due to the whitening transformation and the high number of available pixels within the mini-batch.
@@
Table H1: Cosine similarity between positive pairs (+) and negative pairs (-) during pre-training. P represents a pixel embedding, W represents a waveform embedding. Positive Pairs consist of a pixel embedding and its corresponding pixel/waveform embedding, while negative pairs are computed between a pixel embedding and other pixel/waveform embeddings within the mini-batch.
@@
H.2 Using a Shared Decoder
In our formulation, we used two different decoders for horizontal and vertical structural understanding. We hypothesized that a single real-valued embedding cannot simultaneously encode contrasting information, such as trees of the same species with different vertical structures. Table H2 Shows that in the retrieval case, training a single decoder produces pixel embeddings with no semantic overlap with the waveforms with an average CS of -0.42, while the retrieved pixel embeddings given a pixel input are still highly similar. In the case of training separate decoders, retrieval performance is drastically different, with high correlations (0.99) between pixels and the retrieved waveforms.